OpenStudy (anonymous):

what is the first step to solve the equation 2/x+1/3=4/x A. Add the numerators on the left side. B. Find the LCD. C. Add the denominators on the left side. D. Multiply each side of the equation by 3x^2 In solving the rational equation -1/x-3=2x/x-3 you find that the solutions are -1/2 and 3 however, 3 is an extraneous solution. why? A. There can only be one solution. B. The denominators can't be zero. C. The rational expressions will be zero. D. There was an error in the calculations

OpenStudy (anonymous):

please help! I only have 10 minutes to answer it!

OpenStudy (anonymous):

\[-\frac{4}{x}+\frac{2}{x}+\frac{1}{3}=0 \]Combine the fractions.\[\frac{x-6}{3 x}=0 \]x = 6